Getintoityuh:

The volume of a cone is 392.5 cubic inches. The height of the cone is 15 inches. What is the radius of the cone, rounded to the nearest inch? (Use π = 3.14.) 3 5 13 25

XxXNessalulbaddieXxX:

@smokeybrown

SmokeyBrown:

The formula for volume of a cone is: \[Volume = 1/3*h*\pi*r ^{2}\] Where "r" is the radius, and "h" is the height. Using that formula and the information you're given about the cone, can you solve for "r" and find the radius of the cone?

Getintoityuh:

25?

SmokeyBrown:

25 would be much too large. With that radius, you would get a volume closer to 10000 cubic inches. Remember to use the equation. If you can get "r" by itself on one side, using algebra, then you'll have found your answer

Getintoityuh:

what is it like 5

SmokeyBrown:

It is 5, you're right. We can prove this using the Volume formula 5^2 = 25 25 * 15 = 375 375 * 3.14 = 1177.5 1177.5/3 = 392.5 Good work! And by the way, welcome to QuestionCove!